I have to solve the problem by using substitution Method. If the problem has no solution or an infinite solution or an answer.
1. x + 7Y = 16
-3X + 6Y= 33
2. X + Y = 1
X + Y = 2
This problem has to be sloved by equations by any method.
8x + 9 = -5y
-3x + 3y = -21
In a recent game Will scored 42 points on a combination of 18 two and three point baskets. How many shots of each type were made?

Answer by stanbon(48558)

solve the problem by using substitution Method. If the problem has no solution or an infinite solution or an answer.
1.
1st: x + 7Y = 16
2nd: -3X + 6Y= 33
Solve 1st for x, as follows:
3rd: x=16-7y
Substitute into 2nd to solve for y as follows:
-3(16-7y)+6y=33
-48+21y+6y=33
27y=81
y=3
---------
Substitute in 3rd to solve for x, as follows:
x=16-7(3)
x=-5
Check answer in 2nd as follows:
-3(-5)+6(3)=33
15+18=33
33=33
========================

2.
1st: X + Y = 1
2nd: X + Y = 2
Solve 1st for x, as follows:
x=1-y
Substitute in 2nd to get:
(1-y)+y=2
1=2
This is a contradiction indicating
there is no solution to this system
of equations. It also means you would
get two parallel lines that do not meet
in a point if you were to graph these
two equations.
========================
This problem has to be sloved by equations by any method.
8x + 9 = -5y
-3x + 3y = -21
Hint:
Write both equations in the form ax+b=c
Multiply the 1st by 3 and the 2nd by 8
Then subtract the 1st from the 2nd.
Good luck.
--------------
In a recent game Will scored 42 points on a combination of 18 two and three point baskets. How many shots of each type were made?
Let # of 2-pt baskets be "x"; these are worth 2x points in the game.
The # of 3-pt baskets is "18-x" ; these are worth 3(18-x)=54-3x points.
EQUATION:
points + points = 42
2x+54-3x=42
-x=-12
x=12 (# of 2-point baskets)
18-x=6 (# of 3-point baskets)
Cheers,
Stan H.